# Model Swarms: Collaborative Search to Adapt LLM Experts via Swarm Intelligence

### v1

@@ Abstract
We propose Model Swarms, a collaborative search algorithm to adapt LLMs via swarm intelligence, the collective behavior guiding individual systems. Specifically, Model Swarms starts with a pool of LLM experts and a utility function. Guided by the best-found checkpoints across models, diverse LLM experts collaboratively move in the weight space and optimize a utility function representing model adaptation objectives. Compared to existing model composition approaches, Model Swarms offers tuning-free model adaptation, works in low-data regimes with as few as 200 examples, and does not require assumptions about specific experts in the swarm or how they should be composed. Extensive experiments demonstrate that Model Swarms could flexibly adapt LLM experts to a single task, multi-task domains, reward models, as well as diverse human interests, improving over 12 model composition baselines by up to 21.0% across tasks and contexts. Further analysis reveals that LLM experts discover previously unseen capabilities in initial checkpoints and that Model Swarms enable the weak-to-strong transition of experts through the collaborative search process.
@@ 2 Methodology
$\mathbf{v}_{i}\leftarrow\frac{1}{\mathcal{C}}\big{[}r_{v}\phi_{v}\mathbf{v}_{i}+r_{p}\phi_{p}(\mathbf{p}_{i}-\mathbf{x}_{i})+r_{g}\phi_{g}(\mathbf{g}-\mathbf{x}_{i})-r_{w}\phi_{w}(\mathbf{g}_{w}-\mathbf{x}_{i})\big{]}$
@@ 3 Experiment Settings
Models and Implementation
We implement a prototype of Model Swarms with Gemma-7B (google/gemma-7b-it) (Gemma Team et al., 2024) in the main paper, while we also employ other LLMs such as Mistral-7B (Jiang et al., 2023a) in Table 7. We create a pool of 10 initial experts/particles by fine-tuning Gemma-7B separately on the 10 SFT data domains [^3] We replace the GPT-4 Alpaca subset with Gemini-distilled Alpaca and remove the hardcoded subset. in Tulu-v2 (Ivison et al., 2023) with LoRA (Hu et al., 2022). We fine-tune for 5 epochs with a starting learning rate of 2e-4 and effective batch size of 32 by default. For Model Swarms searches, we employ $N=20$, $\phi_{\lambda}=0.95$, $p=10$, $p_{r}=5$, $\mathcal{K}=50$, while running grid search over other hyperparameters and report the best-found expert based on utility function $f$.
Baselines
We compare with 12 model composition baselines in three categories.
Data and Evaluation
  • Single task: we employ 9 datasets spanning knowledge (MMLU (Hendrycks et al., 2021), MMLU-pro (Wang et al., 2024e), Hellaswag (Zellers et al., 2019)), reasoning (GSM8k (Cobbe et al., 2021), Knowledge Crosswords (Ding et al., 2024), NLGraph (Wang et al., 2024a; Zhang et al., 2024b)), and safety (TruthfulQA (Lin et al., 2022), RealToxicityPrompts (Gehman et al., 2020), AbstainQA (Feng et al., 2024a)). We by default randomly sample 200 and 1000 samples as the validation/test sets: the utility function $f$ is defined as performance on the validation set.
  • Human interest: in addition to preferences represented by reward models, it is crucial to adapt LLM experts directly to human: their preferences, personalized needs, and interest domains. Specifically, 13 human annotators nominated 16 interest domains (e.g., electric vehicles and PhD applications), we then employ Gemini-pro to synthesize 25:25 instructions in each domain as validation/test set. $f$ is defined as LLM-as-a-judge (Zheng et al., 2023) 1-10 scores with Gemini on the validation set, while we evaluate the adaptation to human interest topics on three fronts: improvement in $f$ scores, improvement in factuality with Facts&Evidence (Boonsanong et al., 2024), and human evaluation win rate comparing pre-swarm and post-swarm responses.
4 Results
@@ 5 Analysis
Concretely, the $n$ experts start with a location matrix as an identity matrix $\mathbf{L}=\mathbf{I}_{n\times n}=[\mathbf{l}_{1},\cdots,\mathbf{l}_{n}]$, where the $i$-th row $\mathbf{l}_{i}$ denotes the location of particle $i$: a one-hot vector of 00s and the $i$-th value is 1. For text generation, denoting the next-token probability distribution of expert $i$ as $\mathbf{t}_{i}$, expert $i$’s adjusted token probability becomes $\mathbf{t}^{\prime}_{i}=\sum_{j=1}^{n}\mathbf{l}_{i,j}\mathbf{t}_{j}$ and decode text with $\mathbf{t}^{\prime}_{i}$. In the beginning, $\mathbf{t}^{\prime}_{i}=\mathbf{t}_{i}$ as the expert focuses solely on its own token probabilities. After running updates of location and velocity in the $n$-dimensional search space (Algorithm 1), $\mathbf{t}^{\prime}_{i}$ becomes a composition of $\mathbf{t}$ across experts to optimize $f$. This resembles the collaborative decoding paradigms in existing research (Liu et al., 2024; Shen et al., 2024a), while how to compose the distributions are auto-discovered.
@@
Collaboration of Weak $>$ Strong
@@
Visualizing Search Trajectory
Accelerating with Dropout-K/N
6 Related Work
Composing Diverse LLM Experts
In addition to developing gargantuan general-purpose LLMs, increasing research focus on the composition of multiple models. Mixture-of-experts (MoE) models (Jiang et al., 2024; Lin et al., 2024) and methods (Roller et al., 2021; Lewis et al., 2021; Kudugunta et al., 2021; Pfeiffer et al., 2022; Du et al., 2022; Gururangan et al., 2022; Shen et al., 2024b) are one of the most noted paradigms in composing models, where different approaches vary on parallel neural components (Zhang et al., 2022; Li et al., 2022), routing mechanisms (Zhou et al., 2022; Dai et al., 2022), and expert partition (Gururangan et al., 2023; Jang et al., 2023). More recently, learn-to-fuse approaches propose to “glue” experts together with trainable modules (Bansal et al., 2024), adapters (Wang et al., 2024b), or even LLMs (Jiang et al., 2023b): these approaches often need substantial supervised data and might not be modular to seamlessly add/remove experts. In addition, static model arithmetic approaches propose to compose experts by performing arithmetic on model weights and token probabilities to reconcile sign differences (Yu et al., 2024; Yadav et al., 2024), simulate tuning effects (Liu et al., 2024), and induce compositional capabilities (Ilharco et al., 2023), and more (Davari & Belilovsky, 2023; Jang et al., 2024; Deep et al., 2024; Zheng et al., 2024). In comparison, dynamic model arithmetic proposes to merge models guided by an objective function, employing perplexity heuristics (Mavromatis et al., 2024), evolutionary methods (Akiba et al., 2024), and more (Wortsman et al., 2022a; Huang et al., 2023; Gururangan et al., 2023). Most of these model arithmetic approaches often rely on strong assumptions about the experts how they should be composed (e.g. lion indoors = lion outdoors + (dog indoors - dog outdoors) (Ilharco et al., 2023)). In contrast, Model Swarms presents a modular, assumption-free, and flexible approach to compose and adapt diverse LLM experts guided by as few as 200 data instances.
Evolutionary Algorithms and LLMs
7 Conclusion
Limitations and Ethics Statement
Reproducibility Statement
Acknowledgements
We would like to thank Hanna Mazzawi, Tong Chen, and the UW NLP group for their feedback and Jacob Morrison for the inspiration. We would like to thank Jiajie Yan, Ningnan Wang, Yuhan Liu, Xiaochuang Han, Herun Wan, Leijie Wang, and Wenxuan Ding for their help with human evaluation.
Appendix A Discussion
Three Key Strengths of Model Swarms
Model Swarms and Optimization Research
Non-Neural Reward Models
Long vs. Short
Resilience to Malicious Experts
Model Swarms and Multi-Agent Systems
The role of all “experts” in Model Swarms is homogeneous, i.e. they pursue the same goal/adapt to the same objective as represented by utility function $f$. In multi-agent systems (Rame et al., 2022; Zaman et al., 2023; Ainsworth et al., 2023; Chan et al., 2024; Talebirad & Nadiri, 2023; Chen et al., 2023a; Zhang et al., 2024a; Abdelnabi et al., 2024; Kannan et al., 2023; Zeng et al., 2024; Guo et al., 2024b; Sun et al., 2024; Han et al., 2024; Ishibashi & Nishimura, 2024; Wang et al., 2024d; Zhao et al., 2024; Chen et al., 2024c; Hong et al., 2024; Smit et al., 2024; Chen et al., 2024a; b), the agents often have different roles to jointly complete a task, albeit those roles are more or less hand-crafted and especially through prompting. We envision future work on adapting Model Swarms and automatically discovering heterogeneous and collaborative agents that jointly serve a purpose.
Model Swarms and Model Merging
Appendix B Analysis (cont.)
@@ Appendix A Discussion
$\mathbf{v}_{t}$ $=r_{v}\phi_{v}\mathbf{v}_{t-1}+r_{p}\phi_{p}(\mathbf{p}_{t-1}-\mathbf{x}_{t-1})+r_{g}\phi_{g}(\mathbf{g}_{t-1}-\mathbf{x}_{t-1})-r_{w}\phi_{w}(\mathbf{g}_{w,t-1}-\mathbf{x}_{t-1})$
$=r_{v}\phi_{v}\mathbf{v}_{t-1}+r_{p}\phi_{p}\mathbf{p}_{t-1}-(r_{p}\phi_{p}+r_{g}\phi_{g}-r_{w}\phi_{w})\mathbf{x}_{t-1}+r_{g}\phi_{g}\mathbf{g}_{t-1}-r_{w}\phi_{w}\mathbf{g}_{w,t-1}$
@@
$=\mathbf{x}_{t-1}+\lambda\Big{[}r_{v}\phi_{v}\mathbf{v}_{t-1}+r_{p}\phi_{p}\mathbf{p}_{t-1}-(r_{p}\phi_{p}+r_{g}\phi_{g}-r_{w}\phi_{w})\mathbf{x}_{t-1}+r_{g}\phi_{g}\mathbf{g}_{t-1}-r_{w}\phi_{w}\mathbf{g}_{w,t-1}\Big{]}$
$=\lambda r_{v}\phi_{v}\underline{\mathbf{v}_{t-1}}+\lambda r_{p}\phi_{p}\underline{\mathbf{p}_{t-1}}+\Big{[}1-\lambda(r_{p}\phi_{p}+r_{g}\phi_{g}-r_{w}\phi_{w})\Big{]}\underline{\mathbf{x}_{t-1}}+\lambda r_{g}\phi_{g}\underline{\mathbf{g}_{t-1}}-\lambda r_{w}\phi_{w}\underline{\mathbf{g}_{w,t-1}}$
@@ Appendix C Experiment Details
Figure 15: Search dynamics of per-particle change (left) and global best change (right) of utility function $f$. MMLU-pro, Knowledge Crosswords, Hellaswag, NLGraph, RealToxicityPrompts, and AbstainQA performance are illustrated from top to bottom.

### v2

@@ Abstract
We propose Model Swarms, a collaborative search algorithm to adapt LLMs via swarm intelligence, the collective behavior guiding individual systems. Specifically, Model Swarms starts with a pool of LLM experts and a utility function. Guided by the best-found checkpoints across models, diverse LLM experts collaboratively move in the weight space and optimize a utility function representing model adaptation objectives. Compared to existing model composition approaches, Model Swarms offers tuning-free model adaptation, works in low-data regimes with as few as 200 examples, and does not require assumptions about specific experts in the swarm or how they should be composed. Extensive experiments demonstrate that Model Swarms could flexibly adapt LLM experts to a single task, multi-task domains, reward models, as well as diverse human interests, improving over 12 model composition baselines by up to 21.0% across tasks and contexts. Further analysis reveals that LLM experts discover previously unseen capabilities in initial checkpoints and that Model Swarms enable the weak-to-strong transition of experts through the collaborative search process. Code and data available at https://github.com/BunsenFeng/model_swarm.
@@ 2 Methodology
$\mathbf{v}_{i}\leftarrow\frac{1}{\mathcal{C}}\big{[}r_{v}\phi_{v}\mathbf{v}_{i}+r_{p}\phi_{p}(\mathbf{p}_{i}-\mathbf{x}_{i})$
$+r_{g}\phi_{g}(\mathbf{g}-\mathbf{x}_{i})-r_{w}\phi_{w}(\mathbf{g}_{w}-\mathbf{x}_{i})\big{]}$
@@ 3 Experiment Settings
Models and Implementation
We implement a prototype of Model Swarms with Gemma-7B (google/gemma-7b-it) (Gemma Team et al., 2024) in the main paper, while we also employ other LLMs such as Mistral-7B (Jiang et al., 2023a) in Table 7. We create a pool of 10 initial experts/particles by fine-tuning Gemma-7B separately on the 10 SFT data domains in Tulu-v2 (Ivison et al., 2023) with LoRA (Hu et al., 2022). We fine-tune for 5 epochs with a starting learning rate of 2e-4 and effective batch size of 32 by default. For Model Swarms searches, we employ $N=20$, $\phi_{\lambda}=0.95$, $p=10$, $p_{r}=5$, $\mathcal{K}=50$, while running grid search over other hyperparameters and report the best-found expert based on utility function $f$.
Baselines
We compare with 12 model composition baselines in three categories.
Data and Evaluation
  • Single task: we employ 9 datasets spanning knowledge (MMLU (Hendrycks et al., 2021), MMLU-pro (Wang et al., 2024e), Hellaswag (Zellers et al., 2019)), reasoning (GSM8k (Cobbe et al., 2021), Knowledge Crosswords (Ding et al., 2024), NLGraph (Wang et al., 2024a; Zhang et al., 2024a)), and safety (TruthfulQA (Lin et al., 2022), RealToxicityPrompts (Gehman et al., 2020), AbstainQA (Feng et al., 2024a)). We by default randomly sample 200 and 1000 samples as the validation/test sets: the utility function $f$ is defined as performance on the validation set.
  • Human interest: in addition to preferences represented by reward models, we adapt LLM experts directly to human, their personalized needs and interest domains. Specifically, 13 human annotators nominated 16 interest domains (e.g., electric vehicles and PhD applications), we then employ Gemini-pro to synthesize 25:25 instructions in each domain as validation/test set. $f$ is defined as LLM-as-a-judge (Zheng et al., 2023) 1-10 scores with Gemini on the validation set, while we evaluate the adaptation to human interest topics on three fronts: improvement in $f$ scores, improvement in factuality with Facts&Evidence (Boonsanong et al., 2024), and human evaluation win rate comparing pre-swarm and post-swarm responses.
4 Results
@@ 5 Analysis
Concretely, the $n$ experts start with a location matrix as an identity matrix $\mathbf{L}=\mathbf{I}_{n\times n}=[\mathbf{l}_{1},\cdots,\mathbf{l}_{n}]$, where the $i$-th row $\mathbf{l}_{i}$ denotes the location of particle $i$: a one-hot vector of 00s and the $i$-th value is 1. For text generation, denoting the next-token probability distribution of expert $i$ as $\mathbf{t}_{i}$, expert $i$’s adjusted token probability becomes $\mathbf{t}^{\prime}_{i}=\sum_{j=1}^{n}\mathbf{l}_{i,j}\mathbf{t}_{j}$ and decode text with $\mathbf{t}^{\prime}_{i}$. In the beginning, $\mathbf{t}^{\prime}_{i}=\mathbf{t}_{i}$ as the expert focuses solely on its own token probabilities. After running updates of location and velocity in the $n$-dimensional search space (Algorithm 1), $\mathbf{t}^{\prime}_{i}$ becomes a composition of $\mathbf{t}$ across experts to optimize $f$. This resembles the collaborative decoding paradigms in existing research (Liu et al., 2024; Shen et al., 2024b), while how to compose the distributions are auto-discovered.
@@
6 Related Work
Composing Diverse LLM Experts
In addition to developing gargantuan general-purpose LLMs, increasing research focus on the composition of multiple models. Mixture-of-experts (MoE) models (Jiang et al., 2024; Lin et al., 2024) and methods (Roller et al., 2021; Lewis et al., 2021; Kudugunta et al., 2021; Pfeiffer et al., 2022; Du et al., 2022; Gururangan et al., 2022; Shen et al., 2024a) are one of the most noted paradigms in composing models, where different approaches vary on parallel neural components (Zhang et al., 2022; Li et al., 2022), routing mechanisms (Zhou et al., 2022; Dai et al., 2022), and expert partition (Gururangan et al., 2023; Jang et al., 2023). More recently, learn-to-fuse approaches propose to “glue” experts together with trainable modules (Bansal et al., 2024), adapters (Wang et al., 2024b), or even LLMs (Jiang et al., 2023b): these approaches often need substantial supervised data and might not be modular to seamlessly add/remove experts. In addition, static model arithmetic approaches propose to compose experts by performing arithmetic on model weights and token probabilities to reconcile sign differences (Yu et al., 2024; Yadav et al., 2024), simulate tuning effects (Liu et al., 2024), and induce compositional capabilities (Ilharco et al., 2023), and more (Davari & Belilovsky, 2023; Jang et al., 2024; Deep et al., 2024; Zheng et al., 2024). In comparison, dynamic model arithmetic proposes to merge models guided by an objective function, employing perplexity heuristics (Mavromatis et al., 2024), evolutionary methods (Akiba et al., 2024), and more (Wortsman et al., 2022a; Huang et al., 2023; Gururangan et al., 2023). Most of these model arithmetic approaches often rely on strong assumptions about the experts how they should be composed (e.g. lion indoors = lion outdoors + (dog indoors - dog outdoors) (Ilharco et al., 2023)). In contrast, Model Swarms presents a modular, assumption-free, and flexible approach to compose and adapt diverse LLM experts guided by as few as 200 data instances.
Evolutionary Algorithms and LLMs
7 Conclusion
Impact Statement
Reproducibility Statement
Acknowledgements
We would like to thank Hanna Mazzawi, Tong Chen, and the UW NLP group for their feedback and Jacob Morrison for the inspiration. We would like to thank Jiajie Yan, Ningnan Wang, Yuhan Liu, Xiaochuang Han, Herun Wan, Leijie Wang, and Wenxuan Ding for their help with human evaluation. Shangbin Feng would like to thank the support of the IBM PhD Fellowship and the Jane Street Fellowship.
Appendix A Discussion
Three Key Strengths of Model Swarms
Model Swarms and Optimization Research
Non-Neural Reward Models
Long vs. Short
Resilience to Malicious Experts
Model Swarms and Multi-Agent Systems
The role of all “experts” in Model Swarms is homogeneous, i.e. they pursue the same goal/adapt to the same objective as represented by utility function $f$. In multi-agent systems (Rame et al., 2022; Zaman et al., 2023; Ainsworth et al., 2023; Chan et al., 2024; Talebirad & Nadiri, 2023; Chen et al., 2023a; Zhang et al., 2024b; Abdelnabi et al., 2024; Kannan et al., 2023; Zeng et al., 2024; Guo et al., 2024b; Sun et al., 2024; Han et al., 2024; Ishibashi & Nishimura, 2024; Wang et al., 2024d; Zhao et al., 2024; Chen et al., 2024c; Hong et al., 2024; Smit et al., 2024; Chen et al., 2024a, b), the agents often have different roles to jointly complete a task, albeit those roles are more or less hand-crafted and especially through prompting. We envision future work on adapting Model Swarms and automatically discovering heterogeneous and collaborative agents that jointly serve a purpose.
Model Swarms and Model Merging
Appendix B Analysis (cont.)
Visualizing Search Trajectory
@@ Appendix A Discussion
Accelerating with Dropout-K/N
@@
Collaboration of Weak $>$ Strong
@@
$\mathbf{v}_{t}$ $=r_{v}\phi_{v}\mathbf{v}_{t-1}+r_{p}\phi_{p}(\mathbf{p}_{t-1}-\mathbf{x}_{t-1})+r_{g}\phi_{g}(\mathbf{g}_{t-1}-\mathbf{x}_{t-1})$
$-r_{w}\phi_{w}(\mathbf{g}_{w,t-1}-\mathbf{x}_{t-1})$
$=r_{v}\phi_{v}\mathbf{v}_{t-1}+r_{p}\phi_{p}\mathbf{p}_{t-1}-(r_{p}\phi_{p}+r_{g}\phi_{g}-r_{w}\phi_{w})\mathbf{x}_{t-1}$
$+r_{g}\phi_{g}\mathbf{g}_{t-1}-r_{w}\phi_{w}\mathbf{g}_{w,t-1}$
@@
$=\mathbf{x}_{t-1}+\lambda\Big{[}r_{v}\phi_{v}\mathbf{v}_{t-1}+r_{p}\phi_{p}\mathbf{p}_{t-1}-(r_{p}\phi_{p}+r_{g}\phi_{g}$
$-r_{w}\phi_{w})\mathbf{x}_{t-1}+r_{g}\phi_{g}\mathbf{g}_{t-1}-r_{w}\phi_{w}\mathbf{g}_{w,t-1}\Big{]}$
$=\lambda r_{v}\phi_{v}\underline{\mathbf{v}_{t-1}}+\lambda r_{p}\phi_{p}\underline{\mathbf{p}_{t-1}}+\Big{[}1-\lambda(r_{p}\phi_{p}+r_{g}\phi_{g}$
$-r_{w}\phi_{w})\Big{]}\underline{\mathbf{x}_{t-1}}+\lambda r_{g}\phi_{g}\underline{\mathbf{g}_{t-1}}-\lambda r_{w}\phi_{w}\underline{\mathbf{g}_{w,t-1}}$
@@ Appendix C Experiment Details
Figure 15: Search dynamics of per-particle change (left) and global best change (right) of utility function $f$. MMLU-pro, K-Crosswords, Hellaswag, NLGraph, RealToxicityPrompts, and AbstainQA performance are illustrated from top to bottom.